# Pyramid Vector Quantization for LLMs

### v1

@@ 5 Conclusion
This work explored pyramid vector quantization (PVQ) for quantization of weights and activations in large language models (LLMs). PVQ is a vector quantization method that allows high signal-to-noise ratios without having to build an explicit codebook or perform search. This results in state-of-the-art quantization performance in terms of the most favourable performance to bits-per-weight trade-off, and is amenable to quantization of activation. This has direct practical benefit for post-training model compression, but also opens the door towards quantization at train time. We propose to quantize LLMs using an implicit PVQ codebook on the unit sphere, which can be flexibly configured for codesize and dimensions. On top, we propose a theoretically and empirically motivated way to also quantize gains also enabling small groupsizes in practice. Lastly, we incorporate Hessian information throughout the process to minimize feature error due to quantization. This yields a novel and highly parallelisable algorithm for LLM weights and activations. We demonstrate state-of-the-art quantization performance in terms of superior performance after quantizing pre-trained models, on both weight-only and weight and activation quantization.
@@ Appendix A Mathematical details
A.1 Proof that gains are Beta distributed
Assume we have weights ( $w_{1},w_{2},\ldots,w_{N}$) that are normally distributed: $w_{i}\sim\mathcal{N}(0,\sigma^{2})$. The sum of the squares of these weights for a subset of size ( D ) is chi-squared distributed $u=w_{1}^{2}+w_{2}^{2}+\ldots+w_{G}^{2}\sim\sigma^{2}\chi^{2}_{B}$. Now, consider two independent chi-squared distributed variables: $a\sim\sigma^{2}\chi^{2}_{A},b\sim\sigma^{2}\chi^{2}_{B}$. It can be shown that the ratio of these two variables follows a Beta distribution: $\frac{a}{a+b}\sim\text{Beta}\left(\frac{A}{2},\frac{B}{2}\right)$. This result is independent of the scale parameter ( $\sigma^{2}$ ). To normalize the weights in a group, we consider the ratio: $\frac{w_{1}^{2}+w_{2}^{2}+\ldots+w_{D}^{2}}{w_{1}2+w_{2}^{2}+\ldots+w_{N}^{2}}$, where $D<N$. This can be rewritten as: $\frac{w_{1}^{2}+w_{2}^{2}+\ldots+w_{D}^{2}}{(w_{1}2+w_{2}^{2}+\ldots+w_{D}^{2})+(w_{D+1}^{2}+\ldots+w_{N}^{2})}$. Given that $(w_{1}^{2}+w_{2}^{2}+\ldots+w_{D}^{2})\sim\sigma^{2}\chi^{2}_{D}$ and $(w_{D+1}^{2}+\ldots+w_{D}^{2})\sim\sigma^{2}\chi^{2}_{N-D}$, the ratio follows a Beta distribution: $\frac{w_{1}^{2}+w_{2}^{2}+\ldots+w_{B}^{2}}{w_{1}2+w_{2}^{2}+\ldots+w_{N}^{2}}\sim\text{Beta}\left(\frac{D}{2},\frac{N-D}{2}\right)$. Thus, the normalized weights in a group follow a Beta distribution.
@@ Appendix C Implementation details
C.1 Parallelization using CUDA kernels
@@ C.2 Gain quantization
Appendix D Empirical gain histograms
@@ Appendix E Additional results
E.1 Additional weight-only experiments (shape only)
@@
E.2 Additional quantizing shape and gain experiments
Figure 10: Shape and gain quantization. Test perplexities (PPL) with different quantization methods at various bit levels for gains (within connected sets) and for shape bits (between connected sets).
Figure 11: Shape and gain quantization. Average accuracies (Avg. Acc.) after quantizing with different methods at different bits for gains (within connected set) and various shape bits (between connected sets).
@@ E.4 Additional Zero-shot downstream task experiments
PVQ [3.0 bit shapes, 4 bit gains] 16 3.25 6.85 0.79 0.72 0.74 0.77 0.54 0.60 0.69
PVQ [3.5 bit shapes, 6 bit gains] 16 3.88 6.33 0.81 0.72 0.76 0.80 0.57 0.64 0.72
@@
PVQ [3.0 bit shapes, 4 bit gains] 16 3.25 4.22 0.83 0.76 0.83 0.81 0.57 0.78 0.76
PVQ [3.5 bit shapes, 6 bit gains] 16 3.88 3.99 0.84 0.76 0.84 0.82 0.60 0.78 0.77
@@
PVQ [3.0 bit shapes, 4 bit gains] 16 3.25 7.14 0.78 0.72 0.76 0.74 0.49 0.74 0.70
PVQ [3.5 bit shapes, 6 bit gains] 16 3.88 6.53 0.80 0.73 0.77 0.76 0.51 0.75 0.72

### v2

@@ 5 Conclusion
This work explored pyramid vector quantization (PVQ) for quantization of weights and activations in large language models (LLMs). PVQ is a vector quantization method that allows high signal-to-noise ratios without having to build an explicit codebook or perform search. This results in state-of-the-art quantization performance in terms of the most favourable performance to bits-per-weight trade-off, and is amenable to quantization of activation. This has direct practical benefit for post-training model compression, but also opens the door towards quantization at train time. We propose to quantize LLMs using an implicit PVQ codebook on the unit sphere, which can be flexibly configured for codesize and dimensions. In addition, we propose a theoretically and empirically motivated way to also quantize amplitudes enabling small groupsizes in practice. Lastly, we incorporate Hessian information throughout the process to minimize feature error due to quantization. This yields a novel and highly parallelisable algorithm for LLM weights and activations. We demonstrate state-of-the-art quantization performance in terms of superior performance after quantizing pre-trained models, on both weight-only and weight and activation quantization.
@@ Appendix A Mathematical details
A.1 Proof that amplitudes are Beta distributed
Assume we have weights ( $w_{1},w_{2},\ldots,w_{N}$) that are normally distributed: $w_{i}\sim\mathcal{N}(0,\sigma^{2})$. The sum of the squares of these weights for a subset of size ( D ) is chi-squared distributed $u=w_{1}^{2}+w_{2}^{2}+\ldots+w_{G}^{2}\sim\sigma^{2}\chi^{2}_{B}$. Now, consider two independent chi-squared distributed variables: $a\sim\sigma^{2}\chi^{2}_{A},b\sim\sigma^{2}\chi^{2}_{B}$. It can be shown that the ratio of these two variables follows a Beta distribution: $\frac{a}{a+b}\sim\text{Beta}\left(\frac{A}{2},\frac{B}{2}\right)$ (Frankl & Maehara, 1990). This result is independent of the scale parameter ( $\sigma^{2}$ ). To normalize the weights in a group, we consider the ratio: $\frac{w_{1}^{2}+w_{2}^{2}+\ldots+w_{D}^{2}}{w_{1}2+w_{2}^{2}+\ldots+w_{N}^{2}}$, where $D<N$. This can be rewritten as: $\frac{w_{1}^{2}+w_{2}^{2}+\ldots+w_{D}^{2}}{(w_{1}2+w_{2}^{2}+\ldots+w_{D}^{2})+(w_{D+1}^{2}+\ldots+w_{N}^{2})}$. Given that $(w_{1}^{2}+w_{2}^{2}+\ldots+w_{D}^{2})\sim\sigma^{2}\chi^{2}_{D}$ and $(w_{D+1}^{2}+\ldots+w_{D}^{2})\sim\sigma^{2}\chi^{2}_{N-D}$, the ratio follows a Beta distribution: $\frac{w_{1}^{2}+w_{2}^{2}+\ldots+w_{B}^{2}}{w_{1}2+w_{2}^{2}+\ldots+w_{N}^{2}}\sim\text{Beta}\left(\frac{D}{2},\frac{N-D}{2}\right)$. Thus, the normalized weights in a group follow a Beta distribution.
@@ Appendix C Implementation details
C.1 Dataset
We ran all methods on exactly the same data to ensure fair comparison. We follow the Quarot paper (Ashkboos et al., 2024), and use the same 128 samples of the WikiText-2 dataset and hold-out validation data in all experiments.
C.2 Parallelization using CUDA kernels
@@
Appendix D Empirical amplitude histograms
@@ Appendix E Additional results
E.1 Additional weight-only experiments (direction only)
@@
E.2 Additional quantizing direction and amplitude experiments
Figure 10: direction and amplitude quantization. Test perplexities (PPL) with different quantization methods at various bit levels for amplitudes (within connected sets) and for direction bits (between connected sets).
Figure 11: direction and amplitude quantization. Average accuracies (Avg. Acc.) after quantizing with different methods at different bits for amplitudes (within connected set) and various direction bits (between connected sets).
@@ E.4 Additional Zero-shot downstream task experiments
PVQ [3.0 bit directions, 4 bit amplitudes] 16 3.25 6.85 0.79 0.72 0.74 0.77 0.54 0.60 0.69
PVQ [3.5 bit directions, 6 bit amplitudes] 16 3.88 6.33 0.81 0.72 0.76 0.80 0.57 0.64 0.72
@@
PVQ [3.0 bit directions, 4 bit amplitudes] 16 3.25 4.22 0.83 0.76 0.83 0.81 0.57 0.78 0.76
PVQ [3.5 bit directions, 6 bit amplitudes] 16 3.88 3.99 0.84 0.76 0.84 0.82 0.60 0.78 0.77
@@
PVQ [3.0 bit directions, 4 bit amplitudes] 16 3.25 7.14 0.78 0.72 0.76 0.74 0.49 0.74 0.70
PVQ [3.5 bit directions, 6 bit amplitudes] 16 3.88 6.53 0.80 0.73 0.77 0.76 0.51 0.75 0.72